why will f(x)= x/x^3-2x^2+5x have no zeros? ( THE SLASH IS A FRACTION SYMBOL, EVERYTHING BEFORE THE SLASH IS the nominator AND ANYTHING after the slash is the denominator. )


1
Expert's answer
2022-02-23T14:27:02-0500

Solution:

"f(x)= \dfrac{x}{x^3-2x^2+5x}
\\ \Rightarrow f(x)= \dfrac{x}{x(x^2-2x+5)}
\\ \Rightarrow f(x)= \dfrac{1}{x^2-2x+5}"

To have zeroes, f(x) must be put equal to 0.

"f(x)= 0
\\\Rightarrow \dfrac{1}{x^2-2x+5}=0"

which cannot give any solution as LHS can never be 0.

Thus, no solution.